If k >0 and 25x^2 + bx + 16 = (5x - k)^2 , for all values of x, what is the value of k – b?
(a) -44 (b) -36 (c) 14 (d) 36 (e) 44

Answer by galactus(183)

Since the constant in the given quadratic is 16, then k must equal 4.
If k = 4, then we have (5x-4)^2=25x^2-40x+16 and b=-40
Thus, k-b=4-(-40)=44